OpenStudy (anonymous):

the ones with ( ) will show you the answer I choose :) Simplify (15x2 – 24x + 9) ÷ (3x – 3) = ? ( A. 5x – 3 ) B. 5x + 13 with a –30 remainder C. 5x – 13 with a –30 remainder D. 5x + 3 Simplify (2x + 3)(x – 4) = ? A. 2x2 + 5x + 12 (B. 2x2 + 5x – 12) C. 2x2 – 5x – 12 D. 2x2 – 5x + 12 What is the result if you divide 18r4s5t6= ? -3r2st3 A. –6r2s4t3 B. 6r2s4t3 C. –6r2s5t3 D. 6r2s5t3 not sure about this one tho ^^^

OpenStudy (anonymous):

Check again the second one.

OpenStudy (anonymous):

The first one is good. The last one is the same as the first one but instead of having x's you have r's

OpenStudy (michele_laino):

hint: first question: we have to find the roots of the polynomial 15x^2-24x+9. Do you know how to compute those roots?

OpenStudy (michele_laino):

or, in other words we have to solve this equation: 15x^2-24x+9=0

OpenStudy (anonymous):

no im not smart in algebra never was lol so its why im trying but cant seem to find the answer also sorry for the late reply my wifi is acting up

OpenStudy (michele_laino):

the roots of that equation are given by the subsequent formula: \[\Large \begin{gathered} x = \frac{{12 + \sqrt {{{12}^2} - 15 \times 9} }}{{15}} = ...? \hfill \\ \hfill \\ x = \frac{{12 - \sqrt {{{12}^2} - 15 \times 9} }}{{15}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

can you complete the computations above?

OpenStudy (anonymous):

honestly, no :(

OpenStudy (michele_laino):

here are more steps: \[\Large \begin{gathered} x = \frac{{12 + \sqrt {{{12}^2} - 15 \times 9} }}{{15}} = \frac{{12 + \sqrt {144 - 135} }}{{15}} = \hfill \\ = \frac{{12 + \sqrt 9 }}{{15}} = \frac{{12 + 3}}{{15}} = \frac{{15}}{{15}} = 1 \hfill \\ \end{gathered} \] that is the first root

OpenStudy (michele_laino):

then, the second root is: \[\Large \begin{gathered} x = \frac{{12 - \sqrt {{{12}^2} - 15 \times 9} }}{{15}} = .\frac{{12 - \sqrt {144 - 135} }}{{15}} = \hfill \\ = \frac{{12 - \sqrt 9 }}{{15}} = \frac{{12 - 3}}{{15}} = \frac{9}{{15}} = \frac{3}{5} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so the requested roots are: \[\begin{gathered} x = 1 \hfill \\ x = \frac{3}{5} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, there is a theorem which states that we can rewrite the polynomial 15x^2-24x+9, as below: \[\Large 15{x^2} - 24x + 9 = 5\left( {x - \frac{3}{5}} \right)3\left( {x - 1} \right)\]

OpenStudy (michele_laino):

so your fraction can be rewritten as below: \[\Large \frac{{15{x^2} - 24x + 9}}{{3x - 3}} = \frac{{5\left( {x - \frac{3}{5}} \right)3\left( {x - 1} \right)}}{{3\left( {x - 1} \right)}} = ...?\]

OpenStudy (michele_laino):

please simplify the right side, what do you get?

OpenStudy (anonymous):

when I simplify will I divide or multi?

OpenStudy (michele_laino):

hint: 

OpenStudy (anonymous):

ohh okay so is that what I do with my last question I have up there?

OpenStudy (michele_laino):

you have to cancel similar terms, as I have made above

OpenStudy (michele_laino):

yes! The result is: \[\Large \begin{gathered} \frac{{15{x^2} - 24x + 9}}{{3x - 3}} = \frac{{5\left( {x - \frac{3}{5}} \right)3\left( {x - 1} \right)}}{{3\left( {x - 1} \right)}} = \hfill \\ \hfill \\ = 5\left( {x - \frac{3}{5}} \right) = 5x - 3 \hfill \\ \end{gathered} \] so what is the right option?

OpenStudy (anonymous):

Simplify (2x + 3)(x – 4) = ? C. 2x2 – 5x – 12 because 2x*x + 2x*(-4) + 3*x + 3*(-4) 2x^2 -8x + 3x - 12 2x^2 - 5x - 12 am I correct?

OpenStudy (michele_laino):

your answer is correct!

OpenStudy (anonymous):

see for the last question after the one I just got correct can you please make a picture of that like you didfor those ? idk if you know what im talking about

OpenStudy (michele_laino):

for your last question, we can write: \[\Large \frac{{18{r^4}{s^5}{t^6}}}{{ - 3{r^2}s{t^3}}} = - \frac{{18}}{3} \times {r^{4 - 2}} \times {s^{5 - 1}} \times {t^{6 - 3}} = ...?\]

OpenStudy (michele_laino):

where I have applied the property of division of powers with same bases

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

so its c?

OpenStudy (anonymous):

The process seems x2 more complicate than a long division o.o

OpenStudy (anonymous):